OpenStudy (lisa123):

consider the following function. f(x)=x^2-8, x less than or equal to 0. Find the inverse of f(x)

OpenStudy (anonymous):

let f(x)=y y=x^2-8 x^2=y+8 \[x=-\sqrt{y+8},\] interchange x and y \[y=-\sqrt{x+8},or f ^{-1}\left( x \right)=-\sqrt{x+8}\]

OpenStudy (lisa123):

why is x= -sqrt (y+8). Negative? @surjithayer